This is a Monitoring program using FIES.
Proposal |
Date |
Time |
Type |
Targets |
Files |
Comments |
Status |
|
71-501 |
2025-04-03 |
00:00 |
Monitoring |
- |
- |
No observations due to weather. |
Failed |
|
71-501 |
2025-04-04 |
00:35 |
Monitoring |
EPIC249200379, EPIC249570007 |
FIId040053 - FIId040058 |
Clear skies and seeing around 1.0'' |
Observed |
|
71-501 |
2025-04-08 |
01:00 |
Monitoring |
KIC4260884 |
FIId080044 - FIId080046 |
After multiple attempts to observe with clear skies, OB was completed with clear skies and good seeing. |
Observed |
|
71-501 |
2025-04-09 |
00:00 |
Monitoring |
|
|
100% humidity throughout the night. Could not open. |
Failed |
|
71-501 |
2025-04-10 |
00:00 |
Monitoring |
none |
none |
Not possible to open the dome due to humidity. |
Failed |
|
71-501 |
2025-04-13 |
00:32 |
Monitoring |
EPIC249570007, KIC9652971 |
FIId130102 - FIId130104, FIId130115 - FIId130117 |
|
Observed |
|
71-501 |
2025-04-17 |
00:20 |
Monitoring |
EPIC228720824, EPIC-230748783 |
FIId170033 - FIId170035, FIId170039 - FIId170041 |
Clear skies, one target with seeing 1.0'' and the other 1.4'' |
Observed |
|
71-501 |
2025-04-22 |
00:56 |
Monitoring |
KIC4260884 |
FIId220042 - FIId220044 |
Clear skies and seeing around 1.0'' |
Observed |
|
71-501 |
2025-04-26 |
00:16 |
Monitoring |
EPIC249570007 |
FIId260067 - FIId260069 |
Dark, clear sky, good seeing. |
Observed |
|
71-501 |
2025-04-28 |
00:16 |
Monitoring |
EPIC249200379 |
FIId280094 - FIId280096 |
Clear skies, and the seeing was around 1.5". |
Observed |
|
71-501 |
2025-04-29 |
00:00 |
Monitoring |
- |
- |
The target wasn't observed due to high humidity. The telescope was closed for the night due to
freezing conditions. |
Failed |
|
71-501 |
2025-05-02 |
00:10 |
Monitoring |
EPIC228720824 |
FIIe020065 - FIIe020067 |
Observed with clear skies |
Observed |
|
71-501 |
2025-05-03 |
00:15 |
Monitoring |
EPIC249570007 |
FIIe030100 - FIIe030102 |
Observed with clear skies |
Observed |
|
71-501 |
2025-05-04 |
00:10 |
Monitoring |
EPIC212617037 |
FIIe040113 - FIIe040115 |
Observed with partial thin cloud coverage. |
Observed |
|
71-501 |
2025-05-08 |
00:00 |
Monitoring |
None |
None |
No targets observed due to thick clouds rolling in at 02:30 and we were unable to open
the telescope after that. |
Failed |
|
71-501 |
2025-05-12 |
01:30 |
Monitoring |
EPIC249200379, EPIC249570007, KIC-10454113, KIC-4914923, EPIC-236224056 |
FIIe120085 - FIIe120093, FIIe120096 - FIIe120098, FIIe120109 - FIIe120111 |
Clear skies, and the seeing was oscillating between 1.2" and 1.5".
Autoguiding was not possible only for the targets, EPIC249200379 and EPIC249570007 as they were too close to the moon. |
Observed |
|
Should any of the observations have been carried out using the OB system, you will find additional information
HERE